each of the equal angles of an isosceles triangle measures 25 less than twice the vertex angle. find the measure of the vertex.
2(2x-25)+x=180
4x-50+x=180
5x-50=180
+50 +50
5x=230
/5 /5
x=46
2(46)-25=67 does not equal 180 when put back into equation

Answer by mananth(16946)

2(2x-25)+x=180
4x-50+x=180
5x-50=180
+50 +50
5x=230
/5 /5
x=46
2(46)-25=67
Actually it should be 2*(2(46)-25)+46 =180
..